Question 1 Joshua graduated from college and began working in the family restaurant business. At the end of the third month of the first year he began putting R7440,00 per quarter in an individual retirement account and contributed to it for a total of ten years. The account earned interest at 11% per annum, compounded quarterly. The amount that was available to him after the ten years is
1
Expert's answer
2021-08-19T15:18:02-0400

payment per quarter (P) =R 7440

interest rate per quarter (I) "=11\%/4=2.75\%"

no. of years =10

no. of times compounding per year=4

"N=10\times 4=40"

future value ="\frac{P\times ((1+I)^N-1)}{I}"

Future value"=\frac{7440((1+0.0275)^{40}-1))}{0.0275}=\frac{7440(2.959874-1)}{0.0275}=R530235" (amount available to him after 10 years)